# Here Be Dragons: Bimodal posteriors arise from numerical integration error in longitudinal models

### v1

@@ Abstract
Longitudinal models with dynamics governed by differential equations may require numerical integration alongside parameter estimation. We have identified a situation where the numerical integration introduces error in such a way that it becomes a novel source of non-uniqueness in estimation. We obtain two very different sets of parameters, one of which is a good estimate of the true values and the other a very poor one. The two estimates have forward numerical projections statistically indistinguishable from each other because of numerical error. In such cases, the posterior distribution for parameters is bimodal, with a dominant mode closer to the true parameter value, and a second cluster around the errant value. We demonstrate that bimodality exists both theoretically and empirically for an affine first order differential equation, that a simulation workflow can test for evidence of the issue more generally, and that Markov Chain Monte Carlo sampling with a suitable solution can avoid bimodality. The issue of bimodal posteriors arising from numerical error has consequences for Bayesian inverse methods that rely on numerical integration more broadly.
@@ 1 Introduction
The use of numerical integration in methods for estimating differential equation parameters from observations introduces a source of error (Butcher,, 2016) that is distinct from other processes such as measurement error(Agapiou et al.,, 2014). In the Bayesian inverse method literature, the effect of numerical error on the posterior distribution is under-explored and we wish to present both a case study on problems and a means for investigating them under similar circumstances.
@@
We have identified a specific pathology arising from numerical error: bimodality of posterior parameter distributions for affine first order ODEs. We first observed bimodality arising from numerical methods during the development of the R package hmde (O’Brien et al.,, 2025). The bimodality leads to independent Markov Chain Monte Carlo (MCMC) chains converging to very different, but very specific, parameter combinations: one which was a good estimate of the true values (ie. parameters known from simulation), and a separate very bad estimate. For the erroneous combination, under the chosen numerical method and step size, forward projection produces estimated $Y(t_{j})$ over time that are statistically indistinguishable from the ones produced by the correct parameters, due to numerical error. The posterior is, after all, conditioned on the numerical method used in the model, not the true solution to the integral. As our case study arose from ecological literature, we are thinking about growth functions for DEs, and growth parameters as what is being estimated. The language we use will reflect that background but the situation is more general.
@@
Non-identifiability is a general statistical concept for a situation where multiple estimate values have an indistinguishable relationship to data (Rothenberg,, 1971). The form of non-identifiability we are dealing with is different to cases described in the literature. For example Gelfand and Sahu, (1999) discusses situations where the set of observations is small relative to the parameter space such as can arise in generalised linear mixed effects models. Auger-Méthé et al., (2016) addresses large measurement error as a potential source, and Cockayne et al., (2019) discusses a situation where the posterior distribution is bimodal because there are two true solutions to the underlying function. Numerical error as a source of estimation problems is rarely addressed in the existing literature. The only mentions we found were Agapiou et al., (2014) looking at convergence for MCMC at different levels of discretisation, Cotter et al., (2010) where the step size of an Euler method was found to bias estimates. Posterior multimodality arising from numerical methods appears to be completely novel.
@@
When bimodality arises from numerical error, stability of the integration method matters for potentially a very large region of the parameter space – far beyond true or even plausible parameter combinations – the boundaries of which are probably unknown when first going in. The numerical methods we use (Runge-Kutta 4 and 4-5 order) are extremely common (e.g. Falster et al.,, 2017), and numerically better than other commonly used methods such as Euler (implemented in Cotter et al.,, 2010 and Iida et al.,, 2014). We have identified that the problem arises theoretically, and is empirically observable, for an affine first-order ordinary differential equation with negative slope, such as the von Bertalanffy (Von Bertalanffy,, 1938) model where we first observed it, and which is our central example in a re-parameterised form. Affine ODEs are common, and the RK4 method is typically considered reliable for short step sizes (Butcher,, 2016). To find a situation where the combination of the two produced a novel pathology was unexpected. While the affine model has an analytic solution which can be implemented, that is not always the case.
@@
After demonstrating the mathematical presence of multiple solutions to parameter estimation, we present a simulation workflow that allows the user to test the parameter space for multimodal pathologies before embarking on large-scale analysis. Our testing starts with MCMC estimation, with different numerical methods, priors, and parameter combinations which still produce bimodal posteriors. The problem persists for deterministic estimation methods that use gradient ascent to maximise the likelihood, such as the L-BFGS algorithm (Nocedal and Wright,, 2006). We report the results of experiments to demonstrate how the problem appears in practice, and investigate whether common approaches to dealing with such issues work. We have another function of interest – the extremely non-linear Canham function (Canham et al.,, 2004) – so we use the simulation workflow to test whether we find evidence of bimodal posteriors for parameters in that case as well.
@@
Our use of simulation is distinct from the simulation-based calibration method (Talts et al.,, 2020; Modrák et al.,, 2023) which uses data simulated from the chosen prior to test the consistency of the posterior with rank statistics. We are doing a predominantly qualitative rather than quantitative analysis testing for evidence of the existence of posterior bimodality arising from numerical error, and use simulated data that does not come from the chosen priors. We consider this a reasonable approach because our intent is to demonstrate the existence and mechanism by which numerical bimodality arises, we already have a fix for our use-case in the affine ODE’s analytic solution. In the spirit of Gelman et al., (2020), deliberately breaking a model can be informative, and we aim to present a generally applicable workflow using simulation for testing whether an intended model is susceptible to these novel problems.
1.1 Theory
@@
We will demonstrate the bimodal posterior estimate behaviour for dynamics based on the affine first-order ordinary differential equation (ODE)
$f(Y(t),\beta_{0},\beta_{1})=\beta_{0}-\beta_{1}Y(t),$ (1)
@@
Equation (1) is considered a particularly pathological example for numerical methods (Butcher,, 2016) as it is unbounded above and below, and the negative slope in particular means that negative growth increments can arise from the numerics even if the underlying system is strictly positive. The move from continuous time in the analytic solution to discrete time in the numerical one can result in estimates of $Y(t)$ oscillating around the asymptotic size as the calculated gradient overshoots 0 then is pulled back by being a negative value as seen in Figure 1. Equation (1) serves as an ideal example of the bimodal posterior because of the known pathologies, it is simple with two parameters, and has an analytic solution which allows us to easily access true values for $Y(t)$ without introducing compounding numerical error. The analytic solution is
$Y(t)=\frac{\beta_{0}}{\beta_{1}}+\bigg{(}Y(0)-\frac{\beta_{0}}{\beta_{1}}\bigg{)}\exp(-\beta_{1}t),$ (3)
@@
We have chosen initial conditions $0<Y(0)=1<\frac{\beta_{0}}{\beta_{1}}$, but the asymptote happens regardless as larger sizes with negative slope decrease, and negative sizes with positive slope increase. Equation (3) thus gives a mathematical model that is defined for regions that may not correspond to realistic behaviour in the system being modeled. For example, if $Y$ represents size as in our ecological inspiration, $Y(t)<0$ is physically impossible.
@@
In the posterior bimodality there is a form of non-identifiability (i.e. non-uniqueness of parameters) that arises from numerical errors in a model based on Equation (2). This non-identifiability is separate to other sources of such problems (Latz,, 2023), and currently under-explored in the literature because as far as we can tell it has not been observed before. In this demonstration we use a Runge-Kutta 4th order numerical method (Butcher,, 2016) with different step sizes to show that even ‘small’ step sizes (small enough to have good behaviour at the true parameters) can give problems. A step size in this context is the length of a single iteration of the numerical integral, and is at most the time between observations: $t_{j+1}-t_{j}$. As we are using synthetic data, we have observations 1 unit of time apart, and take our step sizes to be 0.5, 0.25, and 0.125 which divide 1 evenly.
We stress that ‘statistically indistinguishable’ $\hat{Y}(t)$ values in this context means within statistical error of each other, they may not be convergent from an analytic perspective. The statistical equivalence gets involved because we have error and a certain level of resolution in measurement processes. That is, we assume data consists of finite observations of the form $y_{j}$ at time $t_{j}$ that look like
$y_{j}=Y(t_{j})+\text{error},$
and have some finite level of precision. The numerical method may produce estimated values $\hat{Y}(t_{j})$ that differ from the true $Y(t_{j})$ by some amount that is much smaller than the level of precision or observation error for the different parameter combinations, but due to the imprecision of the measurement process the likelihood cannot meaningfully distinguish such $Y(t_{j})$ and $\hat{Y}(t_{j})$. In the Bayesian inverse method literature, this is an issue of practical identifiability, where a finite number of observations combined with measurement error complicate access to the true parameters (Latz,, 2023).
@@
We are attempting to estimate the parameters $\beta_{0}$ and $\beta_{1}$ from observations $y_{j}$, which is based on estimating $\hat{Y}_{j}$ given the prior distribution
@@
and the longitudinal model in Equation (2), which encodes the auto-correlation in the longitudinal data. For this demonstration we are looking at a single simulated individual. To make fitting the model easier we apply a translation by the mean observed size $\bar{y}$, which produces the implemented DE
@@
so the output estimates are consistent with Equation (1). We have default prior distributions for the parameters which are
@@
where the mean and standard deviation for $\beta_{1}$ are for the underlying log-transformed normal distribution. The minimal hierarchical set-up is based on the single individual von Bertalanffy structure from O’Brien et al., (2025), but with a number chosen for the error standard deviation instead of fitting an error parameter.
1.2 What’s going on inside the numerical method?
Before we give a thorough statistical analysis of the posteriors, we demonstrate what the underlying problem is. In the following we compare the numerical behaviour of the Runge-Kutta 4th (RK4) order algorithm with step size 0.5 at the true parameter values we are working with for this demonstration ($\beta_{0}=10$ and $\beta_{1}=1$), and the behaviour of the same algorithm at the ‘bad’ estimate point we found for that method and step size: $\hat{\beta}_{0}=49.3$ and $\hat{\beta}_{1}=4.91$.
@@ 1.2 What’s going on inside the numerical method?
In the case of Equation (1) the time parameter $t$ does not appear explicitly so our $k$ formulas do not depend on it. We will refer to where the $k$ parameters are estimated as ‘sub-steps’. Note that $t_{j}$ and $t_{j+1}=t_{j}+h$ are typically different to the times at which observations occur. The observation interval is an upper bound on the numerical step size $h$, but we can take multiple numerical steps between the observations.
@@
It is precisely because the bad numerical method for $\beta_{0}=49.3$ and $\beta_{1}=4.91$ give $\hat{Y}(t_{j})$ values so close to the analytic solution when $\beta_{0}=10$ and $\beta_{1}=1$ that we see the estimation problems. For a visual comparison see Figure 1(a) and (b). The estimation algorithm will converge to any parameter combination it finds that gives $\hat{Y}(t_{j})$ close to the observations provided. In these cases the algorithm finds a point far from the true value where size estimates match the observations extremely well because of error in the numerical method. Non-identifiability here is not a problem with MCMC or another chosen estimation method, it is a structural issue arising from the use of numerical methods.
@@
We can see the problem in plots of the estimated values for $Y(t)$ and $f(Y(t))$ at each step and sub-step for time from 0 to 4. Figure 1 shows instability of the numerical method at $\beta_{0}=49.3$, $\beta_{1}=4.91$, and stability at $\beta_{0}=10$, $\beta_{1}=1$. Both converge to the same asymptotic size $\beta_{0}/\beta_{1}=10$, but because we are using the numerical method to estimate longitudinal data, we care about the stability at the intermediate times as well. The main feature of interest is the huge variation of the $\hat{Y}$ estimate in substeps given the erroneous parameter values, which comes from the gradient oscillating between positive and negative values in Panels (a) and (c). The excessive size happens when the increment estimate at a step $\hat{Y}(t_{1})$ puts the next value $\hat{Y}(t_{2})$ past the asymptotic limit $\beta_{0}/\beta_{1}$, where the gradient is negative. Because the gradient at $\hat{Y}(t_{2})$ is negative, the subsequent $\hat{Y}(t_{3})$ value will be smaller. The numerical method at the pathological parameters is behaving so badly that we actually see a negative $\hat{Y}$ value in the first integration’s second substep.
Figure 2: Plot of Equation (5) with $Y(t_{j})=1$ and $\alpha=10$ for first five steps of size $h=0.5$. The intersections with 0 are at $\hat{\beta}_{1}=1.0008$ and $\hat{\beta}_{1}=4.9112$.
In the case of Equation (1) we can leverage an explicit from of the RK4 algorithm in Equation (4) to get a polynomial in $\beta_{0}$ and $\beta_{1}$ [^1] Computation in the supplementary material.. We have to take care to differentiate the true parameters used in the analytic solution from the parameters we estimate from the longitudinal model, so we denote the estimated parameters with $\hat{\beta}_{0}$ and $\hat{\beta}_{1}$. For a step size of $h$, given initial size $Y(t_{j})$ and the size at the next numerical step from the analytic solution $Y(t_{j}+h)$ we get
$0=$ $\ Y(t_{j})-Y(t_{j}+h)+h\hat{\beta}_{0}-\hat{\beta}_{1}\bigg{(}hY(t_{j})+\frac{h^{2}}{2}\hat{\beta}_{0}\bigg{)}+\hat{\beta}_{1}^{2}\bigg{(}\frac{h^{2}}{2}Y(t_{j})+\frac{h^{3}}{6}\hat{\beta}_{0}\bigg{)}$
$\ -\hat{\beta}_{1}^{3}\bigg{(}\frac{h^{3}}{6}Y(t_{j})+\frac{h^{4}}{24}\hat{\beta}_{0}\bigg{)}+\hat{\beta}_{1}^{4}\frac{h^{4}}{24}Y(t_{j}).$

### v2

@@ Abstract
Longitudinal models with dynamics governed by differential equations may require numerical integration alongside parameter estimation. We have identified a situation where the numerical integration introduces error in such a way that it becomes a novel source of non-uniqueness in estimation. We obtain two very different sets of parameters, one of which is a good estimate of the true values and the other a very poor one. The two estimates have forward numerical projections statistically indistinguishable from each other because of numerical error. In such cases, the posterior distribution for parameters is bimodal, with a dominant mode closer to the true parameter value, and a second cluster around the errant value. We demonstrate that multi-modality exists both theoretically and empirically for an affine first order differential equation, that a simulation workflow can test for evidence of the issue more generally, and that Markov Chain Monte Carlo sampling with a suitable solution can avoid bimodality. The issue of multi-modal posteriors arising from numerical error has consequences for Bayesian inverse methods that rely on numerical integration more broadly.
@@ 1 Introduction
The use of numerical integration in methods for estimating differential equation parameters from observations introduces a source of error (Butcher,, 2016) that is distinct from other processes such as measurement error (Agapiou et al.,, 2014). In the Bayesian literature, the effect of numerical error on the posterior distribution is under-explored and we wish to present a case study on identifiability problems which doubles as a procedure for investigating them under similar circumstances. In this paper, we describe a specific pathology arising from numerical error: erroneous multi-modality of posterior parameter distributions for affine first order ODEs where there are multiple parameter combinations that produce the same numerical solution. Such multi-modality is a previously undescribed form of non-identifiability in Bayesian methods.
@@
We first observed bimodality arising from numerical methods while testing a longitudinal model for the von Bertalanffy function (Von Bertalanffy,, 1938). Posterior bimodality leads to independent Markov Chain Monte Carlo (MCMC) chains converging to very different, but very specific, parameter combinations, one which is a good estimate of the true values (i.e. parameters known from simulation), and a separate very bad estimate. For the erroneous combination, under the chosen numerical method and step size, forward projection produces estimates of $Y(t_{j})$ that are the same as those produced by the correct parameters, due to numerical error. The posterior is, after all, conditioned on the numerical method used in the model, not the true solution to the integral. As our case study arose from ecological literature, We are thinking about growth functions for DEs, and growth parameters as what is being estimated. However the example is a simple DE likely to appear in other examples. The language we use will reflect the ecological background but the situation is more general.
@@
Non-identifiability is a general statistical concept for a situation where multiple estimate values have an indistinguishable relationship to data (Rothenberg,, 1971). The form of non-identifiability We are dealing with is different to cases described in the literature. For example Gelfand and Sahu, (1999) discusses situations where the set of observations is small relative to the parameter space such as can arise in generalised linear mixed effects models. Auger-Méthé et al., (2016) addresses large measurement error as a potential source, and Cockayne et al., (2019) discusses a situation where the posterior distribution is bimodal because there are two true solutions to the underlying function. Numerical integration error as a source of estimation problems is rarely addressed in the existing literature. The only mentions we found were Agapiou et al., (2014) looking at convergence for MCMC at different levels of discretisation and Cotter et al., (2010) where the step size of an Euler method was found to bias estimates. In supplementary material to O’Brien et al., (2024) we investigated bias arising from numerical error for a linear ODE. Posterior multi-modality arising from numerical methods appears to be completely novel.
@@
When multi-modality arises from numerical error, stability of the integration method matters for potentially a very large region of the parameter space – far beyond true or even plausible parameter combinations – the boundaries of which are probably unknown when first going in. The numerical methods we use (Runge-Kutta 4 and 4-5 order) are extremely common (e.g. Falster et al.,, 2017), and numerically more reliable than other commonly used methods such as Euler (implemented in Cotter et al.,, 2010 and Iida et al.,, 2014). We demonstrate that the problem arises theoretically, and is empirically observable, for an affine first-order ordinary differential equation with negative slope, such as the von Bertalanffy (Von Bertalanffy,, 1938) model where we first observed it. The von Bertalanffy model is our central example in a re-parameterised form. Affine ODEs are common, and the RK4 method is typically considered reliable for short step sizes (Butcher,, 2016). To find a situation where the combination of the two produced a novel pathology was unexpected. The affine ODE has an analytic solution which allows usto side-step the identifiability problem by not using numerical methods, however if numerical error introduces non-uniqueness elsewhere it may be in situations without such a work-around.
@@
In this paper we first prove that in general, affine ODEs subject to RK4 numerical integration in a longitudinal model will have multi-modal posteriors. We then present a simulation workflow for matching empirical results that allows a user to test the parameter space of more general models for multi-modal pathologies before embarking on large-scale analysis. Our testing starts with MCMC estimation, with different numerical methods, priors, and parameter combinations which still produce bimodal posteriors. The problem persists for deterministic estimation methods that use gradient ascent to maximise the likelihood, such as the L-BFGS algorithm (Nocedal and Wright,, 2006). We report the results of experiments to demonstrate how the problem appears in practice, and investigate whether common approaches to dealing with such issues work. We have another function of interest – the extremely non-linear Canham function based on a log-normal distribution function from Canham et al., (2004) – so we use the simulation workflow to test whether we find evidence of bimodal posteriors for parameters in that case as well.
@@
Our use of simulation is distinct from a simulation-based calibration method (Talts et al.,, 2020; Modrák et al.,, 2023) which uses data simulated from the chosen prior to test the consistency of the posterior with rank statistics. We are doing a predominantly qualitative rather than quantitative analysis testing for evidence of the existence of posterior bimodality arising from numerical error, and use simulated data that does not come from the chosen priors. We consider this a reasonable approach because our intent is to demonstrate the existence and mechanism by which numerical bimodality arises, we already have a fix in the affine ODE’s analytic solution. In the spirit of Gelman et al., (2020), deliberately breaking a model can be informative, and we aim to present a generally applicable workflow using simulation for testing whether an intended model is susceptible to these novel problems.
2 Theoretical analysis
@@
We demonstrate the bimodal posterior estimate behaviour for dynamics based on the affine first-order ordinary differential equation (ODE)
$f(Y(t),\beta_{0},\beta_{1})=\beta_{0}-\beta_{1}Y(t),$ (2.1)
@@
Equation (2.1) is considered a particularly pathological example for numerical methods (Butcher,, 2016) as it is unbounded above and below, and the negative slope in particular means that negative growth increments can arise from the numerics even if the underlying system is strictly positive. The move from continuous time in the analytic solution to discrete time in the numerical one can result in estimates of $Y(t)$ oscillating around the asymptotic size as the calculated gradient overshoots 0 then is pulled back by being a negative value as seen in Figure 1. Equation (2.1) serves as an ideal example of the bimodal posterior because of the known pathologies, it is simple with two parameters, and has an analytic solution which allows us to easily access true values for $Y(t)$ without introducing compounding numerical error. The analytic solution is
$Y(t)=\frac{\beta_{0}}{\beta_{1}}+\bigg{(}Y(0)-\frac{\beta_{0}}{\beta_{1}}\bigg{)}\exp(-\beta_{1}t),$ (2.3)
@@
We choose the initial condition $0<Y(0)=1<\frac{\beta_{0}}{\beta_{1}}$. Equation (2.3) gives a mathematical model that is defined for regions that may not correspond to realistic behaviour in the system being modelled. For example, if $Y$ represents physical size, $Y(t)<0$ is impossible. Negative growth is also possible if $Y(t)>\frac{\beta_{0}}{\beta_{1}}$.
@@
In this demonstration we use a Runge-Kutta 4th order numerical method (Butcher,, 2016) with different choices for the step size to show that even ‘small’ step sizes (small enough to have good behaviour at the true parameters) can give problems. A step size in this context is the length of a single iteration of the numerical integral, and is at most the time between observations: $t_{j+1}-t_{j}$. As We are using synthetic data, we have observations 1 unit of time apart, and take step sizes to be 0.5, 0.25, and 0.125 which divide 1 evenly.
We assume data consists of finite observations of the form $y_{j}$ at time $t_{j}$ that look like
$y_{j}=Y(t_{j})+\text{error},$ (2.4)
and have some finite level of precision. In the Bayesian literature, this set-up is an issue of practical identifiability, where a finite number of observations combined with measurement error complicate access to the true parameters (Latz,, 2023). We use finite true $Y(t_{j})$ values for theoretical results, and observations with error $y_{j}$ for empirical ones.
@@
We are attempting to estimate the parameters $\beta_{0}$ and $\beta_{1}$ from simulated observations $y_{j}$ of the form in Equation (2.4). We estimate $\hat{Y}_{j}$ given the prior distribution
@@
and the longitudinal model in Equation (2.2), which encodes the auto-correlation in the data. For this demonstration We are looking at a single simulated individual. To make fitting the model easier we apply a translation by the mean observed size $\bar{y}$, which produces the implemented DE
@@
so the output estimates are consistent with Equation (2.1). We have default prior distributions for the parameters which are
@@
where the mean and standard deviation for $\beta_{1}$ are for the underlying log-transformed normal distribution. Sensitivity of estimation to the standard deviation is tested as part of the experiments. The minimal hierarchical set-up is based on a single individual structure such as used in O’Brien et al., (2024), but with a number chosen for the error standard deviation instead of fitting an error parameter.
2.1 Numerical integration theory
In this section we prove non-identifiability of the affine model parameters under a Runge-Kutta 4th order numerical method (RK4) in a theoretical context using the true values rather than observed data. We prove that, amongst the set of solutions for which the asymptotic size is $\alpha$, the affine ODE has multiple $\hat{\beta}_{1}$ parameters that will produce the same sizes over time for the RK4 method, then look at the specific case study where $\beta_{0}=10$ and $\beta_{1}=1$.
@@ 2.1 Numerical integration theory
In the case of Equation (2.1) the time parameter $t$ does not appear explicitly so our $k$ formulas do not depend on it. We will refer to where the $k$ parameters are estimated as ‘sub-steps’. Note that $t_{j}$ and $t_{j+1}=t_{j}+h$ are typically different to the times at which observations occur. The observation interval is an upper bound on the numerical step size $h$, but we can take multiple numerical steps between the observations.
@@
In a chosen situation where $\beta_{0}=10$ and $\beta_{1}=1$ which produces a bimodal posterior, at the second (erroneous) parameter combination the numerical method is sufficiently unstable that forward projection produces $\hat{Y}(t_{j})$ values that match the analytic solution at the true parameters. For a visual comparison see Figure 1(a) and (b), which leverages the bad estimates ($\beta_{0}=49.3,\beta_{1}=4.91$) paired to our chosen parameters $\beta_{0}=10$ and $\beta_{1}=1$. The estimation algorithm will converge to any parameter combination that gives $\hat{Y}(t_{j})$ close to the observations provided. In these cases the algorithm finds a point far from the true value where size estimates match the observations extremely well because of numerical integration error. Non-identifiability here is not a problem with MCMC or another chosen estimation method, it is a structural issue arising from the use of numerical integration in the longitudinal structure.
@@
We can see the problem in plots of the estimated values for $Y(t)$ and $f(Y(t))$ at each step and sub-step for time from 0 to 4. Figure 1 shows instability of the numerical method at $\beta_{0}=49.3$, $\beta_{1}=4.91$ (the bad parameter combination for this example we found both empirically and theoretically), and stability at $\beta_{0}=10$, $\beta_{1}=1$. Both parameter combinations converge to approximately the same asymptotic size $\beta_{0}/\beta_{1}=10$, so in a sense the long-term numerics are stable, but because We are using the numerical method to fit longitudinal data, we care about the stability at the intermediate times as well. The main feature of interest is the huge variation of the $\hat{Y}$ estimate in sub-steps given the erroneous parameter values, which comes from the gradient oscillating between positive and negative values in Panels (a) and (c). The excessive size happens when the increment estimate at a step $\hat{Y}(t_{1})$ puts the next value $\hat{Y}(t_{2})$ past the asymptotic limit $\beta_{0}/\beta_{1}$, where the gradient is negative. Because the gradient at $\hat{Y}(t_{2})$ is negative, the subsequent $\hat{Y}(t_{3})$ value will be smaller. The numerical method at the pathological parameters is behaving so badly that we actually see a negative $\hat{Y}$ value in the first integration’s second sub-step.
The following provides proofs for non-identifiability in general for affine ODEs when using the RK4 numerical method. In the case of the affine ODE we can leverage an explicit from of the RK4 algorithm in Equation (2.5) to get a polynomial in $\beta_{0}$ and $\beta_{1}$. We prove non-identifiability for solutions that satisfy:
$\frac{\beta_{0}}{\beta_{1}}=\alpha,$
which ensures that the asymptotic size is $\alpha$. Our numerical work suggests that in practice, the asymptotic size typically has a strong signal in data, and estimates of the parameters – $\hat{\beta}_{0}$ and $\hat{\beta}_{1}$ – tend to reflect this, with $\hat{\beta}_{0}/\hat{\beta}_{1}\approx\alpha$. By using the transformation $\beta_{0}=\alpha\beta_{1}$ we derive a polynomial in $\beta_{1}$ for the numerical step in the following lemma.
Lemma 2.1.
For a first order affine ODE of the form
$\frac{dY}{dt}=\beta_{0}-\beta_{1}Y(t),$
consider $\beta_{0}=\alpha\beta_{1}$ (that is, parameter combinations for which the asymptotic size is $\alpha$). A single step of the Runge-Kutta 4th order numerical integration is given by
$\hat{Y}(t_{j}+h)=Y(t_{j})+(\alpha-Y(t_{j}))\bigg{(}h\beta_{1}-\frac{h^{2}}{2}\beta_{1}^{2}+\frac{h^{3}}{6}\beta_{1}^{3}-\frac{h^{4}}{24}\beta_{1}^{4}\bigg{)}$ (2.6)
where $h$ is the step size and $\alpha=\beta_{0}/\beta_{1}$ is the asymptotic $Y$ value.
Proof.
The Runge-Kutta scheme for $Y(t)$ given $\frac{dY}{dt}=f(Y(t),t,\theta)$ is calculated as
$\hat{Y}(t_{j}+h)=Y(t_{j})+\frac{h}{6}(k_{1}+2k_{2}+2k_{3}+k4)$ (2.7)
where $h$ is the step size and $k$ values are estimates of the gradient at iterated points given by
$k_{1}=$ $\ f(Y(t_{j}),t_{j},\theta),$
$k_{2}=$ $\ f(Y(t_{j})+hk_{1}/2,t_{j}+h/2,\theta),$
$k_{3}=$ $\ f(Y(t_{j})+hk_{2}/2,t_{j}+h/2,\theta),$
$k_{4}=$ $\ f(Y(t_{j})+hk_{3},t_{j}+h,\theta).$
First we calculate formulas for each $k_{i}$ that depend only on $\beta$s and $Y(t_{j})$ by substituting in the function definition sequentially. We start with
$k_{1}=\beta_{0}-\beta_{1}Y(t_{j}),$
which we substitute into $k_{2}$ to get
$k_{2}=$ $\ \beta_{0}-\beta_{1}\bigg{(}Y(t_{j})+\frac{h}{2}k_{1}\big{)}$
$=$ $\ \beta_{0}-\beta_{1}\bigg{(}Y(t_{j})+\frac{h}{2}(\beta_{0}-\beta_{1}Y(t_{j}))\bigg{)}$
$=$ $\ \beta_{0}-\beta_{1}\bigg{(}Y(t_{j})+\frac{h}{2}\beta_{0}\bigg{)}+\beta_{1}^{2}\frac{h}{2}Y(t_{j}).$
Subsequent $k$s are
$k_{3}=$ $\ \beta_{0}-\beta_{1}\bigg{(}Y(t_{j})+\frac{h}{2}\beta_{0}\bigg{)}+\beta_{1}^{2}\bigg{(}\frac{h}{2}Y(t_{j})+\frac{h^{2}}{4}\beta_{0}\bigg{)}-\beta_{1}^{3}\frac{h^{2}}{4}Y(t_{j}),$
$k_{4}=$ $\ \beta_{0}-\beta_{1}(Y(t_{j})+h\beta_{0})+\beta_{1}^{2}\bigg{(}hY(t_{j})+\frac{h^{2}}{2}\beta_{0}\bigg{)}$
$\ -\beta_{1}^{3}\bigg{(}\frac{h^{2}}{2}Y(t_{j})+\frac{h^{3}}{4}\beta_{0}\bigg{)}+\beta_{1}^{4}\frac{h^{4}}{4}Y(t_{j}).$
To extract the polynomial we substitute the $k$s into Equation 2.7 and do some re-organising of terms to get
$\hat{Y}(t_{j}+h)=$ $\ Y(t_{j})+\frac{h}{6}\Bigg{(}6\beta_{0}-\beta_{1}\big{(}6Y(t_{j})+3h\beta_{0}\big{)}+\beta_{1}^{2}\big{(}3hY(t_{j})+h^{2}\beta_{0}\big{)}$
$\qquad\qquad\qquad\quad-\beta_{1}^{3}\bigg{(}h^{2}Y(t_{j})+\frac{h^{3}}{4}\beta_{0}\bigg{)}+\beta_{1}^{4}\frac{h^{3}}{4}Y(t_{j})\Bigg{)}$
$=$ $\ Y(t_{j})+h\beta_{0}-\beta_{1}\bigg{(}hY(t_{j})+\frac{h^{2}}{2}\beta_{0}\bigg{)}+\beta_{1}^{2}\bigg{(}\frac{h^{2}}{2}Y(t_{j})+\frac{h^{3}}{6}\beta_{0}\bigg{)}$
$\qquad\qquad\qquad\quad-\beta_{1}^{3}\bigg{(}\frac{h^{3}}{6}Y(t_{j})+\frac{h^{4}}{24}\beta_{0}\bigg{)}+\beta_{1}^{4}\frac{h^{4}}{24}Y(t_{j}).$